the sqaure root of 2x+5 = the sqaure root of 3x+1
Answer by venugopalramana(3286)

the sqaure root of 2x+5 = the sqaure root of 3x+1...SQUARE BOTH SIDES WE GET
2X+5=3X+1
3X-2X=5-1=4
X=4